1) Because of the sales tax in a particular region, a customer in a supermarket must pay $2.14 for a loaf of bread marked $2.00. What is the effective tax rate on this purchase, expressed as a percentage?

2)A pet lamb grows rapidly, with its mass proportional to the cube of its length. When the lamb's length changes by 15.4%, its mass increases by 15.4 kg. Find the lamb's mass at the end of this process.

3) A thin rod extends from x = 0 to x = 17.0 cm it has a cross-sectional area A = 6.50 cm2, and its density increases uniformly in the positive x-direction from 3.00 g/cm3 at one endpoint to 18.0 g/cm3 at the other. a) The density as a function of distance for the rod is given by ρ = B + Cx, where B and C are constants. What are the values of B (in g/cm3) and C (in g/cm4). b) Finding the total mass of the rod requires integrating the density function over the entire length of the rod.
Answer #1

1) Effective Tax in amount is (2.14 - 2.00)$= 0.14$

Tax rate in percentage = 0.14/2.00 * 100 = 7%

2)